OpenStudy (anonymous):

-2 -( -3)-4 -(6)+(-7)

OpenStudy (anonymous):

-2+3-4-6+9+(-7)= -7

OpenStudy (anonymous):

Do you understand how I did that? I was working from left to right... work with one set of numbers at a time till you are left with the final answer

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-2+3= 1 So now you are left with 1-4-6+9+(-7)= 1-4=-3 so now you have -3-6+9+(-7)= -3-6=-9 so now you have -9+9=0 so now you have 0+(-7)= -7

OpenStudy (anonymous):

thx